OpenStudy (firejay5):

Solve each system of equations by using substitution: Show work and steps on how you got it. 18. 2x + 4y = 6 7x = 4 + 3y

OpenStudy (anonymous):

Hello :)

OpenStudy (firejay5):

Hello?

OpenStudy (anonymous):

so the best way to solve this, is to set the equation to Ax+By = C

OpenStudy (anonymous):

the first one is already there so the second one you want to put the y with the x on the left side of the equation

OpenStudy (anonymous):

follow?

zepdrix (zepdrix):

Woops, are you thinking of elimination Calle? Cause the directions told him to use substitution :D

OpenStudy (anonymous):

Oh, good catchzep

OpenStudy (anonymous):

ok for substitution you want to get one of the equations to say x= By + C or y = Ax + C

OpenStudy (anonymous):

the first one is quite easy if you just solve for x by subtracting 4y and dividing by 2

OpenStudy (anonymous):

2x+4y=6 2x=-4y+6 x=-2y+3

OpenStudy (anonymous):

got that part?

OpenStudy (anonymous):

you there?

OpenStudy (firejay5):

yea, just letting you know this isn't standard form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so plug x into the 2nd equation where x is

OpenStudy (firejay5):

got that part @Calle87

OpenStudy (anonymous):

Fire what is your problem?

OpenStudy (firejay5):

substituting and stuff

OpenStudy (anonymous):

Substitution is basically like taking the definition of one variable and plugging it into an equation. If x = y + 4 and x + y = 9 If x = y + 4, cant you just replace that with the x in the second equation?

OpenStudy (firejay5):

I am substituting x = -2y + 3 into the second equation and I am trying to get y by itself

OpenStudy (anonymous):

Why x = -2y + 3?

OpenStudy (anonymous):

Ah, I see. You solved the first one.. Ok

OpenStudy (firejay5):

cause that's my answer to the first equation

OpenStudy (anonymous):

So whats the problem? dont you get just y??

OpenStudy (firejay5):

this is substitution

OpenStudy (anonymous):

Yes and....

OpenStudy (anonymous):

You are substituting x = -2y + 3 into the second equation

OpenStudy (firejay5):

yes I am wondering if it = y = \[\frac{ 17 }{ 11 }\]

OpenStudy (anonymous):

7x = 4 + 3y 7(-2y + 3) = 4 + 3y -14y + 21 = 4 + 3y Check your work.

OpenStudy (firejay5):

yea that's right then combine like terms -11y + 21 = 4, then subtract 21 from both sides ---> -11y = -17 <--- y = 17/11

OpenStudy (anonymous):

Nope wait.. -14y + 21 = 4 + 3y -3y - 3y ------------------ -17y + 21 = 4 See your error?

OpenStudy (firejay5):

no we combine our like terms

OpenStudy (anonymous):

Yes but the 3y is on the other side of the equal sign, so you must subtract it from both sides.

OpenStudy (firejay5):

so y = 1

OpenStudy (anonymous):

Yep!

OpenStudy (firejay5):

OKAY!

OpenStudy (anonymous):

:) If I helped you, would you mind giving me a medal? Thanks!

OpenStudy (firejay5):

Well until I get (x,y)

OpenStudy (anonymous):

Ok, do you understand how to get x?

OpenStudy (firejay5):

yes! I plug y into first equation. Would I subtract 2x or add 4x first

OpenStudy (anonymous):

Where would you get 4x?

OpenStudy (firejay5):

multiplying -4 (1)

OpenStudy (firejay5):

1 is from y =1

OpenStudy (anonymous):

Ok so it is.. 2x + 4(1) = 6 2x + 4 = 6 - 4 - 4 --------------

OpenStudy (firejay5):

I thought it was 2x = 4x + 6

OpenStudy (anonymous):

I could see where you got that... This is a time where you got to be careful of silly mistakes ;)

OpenStudy (firejay5):

oh okay

OpenStudy (anonymous):

If you need any more help, just ask

OpenStudy (firejay5):

the point is (5,1)

OpenStudy (anonymous):

No it cant be. Try plugging it into the equations, and they wont be equal.

OpenStudy (firejay5):

you know how you get a point. For example (3,-2)

OpenStudy (anonymous):

Through the process we just did? You mean graphing it?

OpenStudy (firejay5):

no not graphing it

OpenStudy (anonymous):

then doing what to it?

OpenStudy (firejay5):

you know how figuring out what x and y =

OpenStudy (firejay5):

of one or both equations

OpenStudy (anonymous):

Oh yes Didnt we just do that through Substitution?

OpenStudy (firejay5):

yea